# Flatness is Necessary, Neural Collapse is Not: Rethinking Generalization via Grokking

### v1

@@ Abstract
Neural collapse, i.e., the emergence of highly symmetric, class-wise clustered representations, is frequently observed in deep networks and is often assumed to reflect or enable generalization. In parallel, flatness of the loss landscape has been theoretically and empirically linked to generalization. Yet, the causal role of either phenomenon remains unclear: Are they prerequisites for generalization, or merely by-products of training dynamics? We disentangle these questions using grokking, a training regime in which memorization precedes generalization, allowing us to temporally separate generalization from training dynamics and we find that while both neural collapse and relative flatness emerge near the onset of generalization, only flatness consistently predicts it. Models encouraged to collapse or prevented from collapsing generalize equally well, whereas models regularized away from flat solutions exhibit delayed generalization. Furthermore, we show theoretically that neural collapse implies relative flatness under classical assumptions, explaining their empirical co-occurrence. Our results support the view that relative flatness is a potentially necessary and more fundamental property for generalization, and demonstrate how grokking can serve as a powerful probe for isolating its geometric underpinnings.
@@ 1 Introduction
In this paper, we challenge the common conjecture that NC is necessary for generalization (Mixon et al., 2022; Zhou et al., 2022; Súkeník et al., 2023; Zhu et al., 2021). Our key idea is to leverage grokking as a unique observational window: since generalization emerges only after prolonged memorization, it enables a clean separation between training dynamics and generalization. This allows us to ask—do neural collapse and flatness merely correlate with generalization, or do they contribute to it?
We measure both phenomena in the penultimate layer. For flatness, we adopt the relative flatness metric proposed by Petzka et al. (2021), which considers the Hessian trace normalized by the weight norm, a quantity theoretically and empirically shown to align with generalization. To make this computable in large state-of-the-art neural networks, we employ the alternative closed-form upper bound introduced by (Walter et al., 2024), which is valid in the penultimate layer under cross-entropy loss. For NC, we measure the class-wise clustering of penultimate-layer representations through empirical variance and mean properties, tracking its emergence throughout training. In grokking experiments, we find that flatness emerges alongside generalization, never before. NC, in contrast, appears earlier during memorization, suggesting that it may result from, rather than enable, generalization (cf. Fig. 1).
@@
To isolate their functional roles, we perturb training dynamics. When we suppress collapse-style clustering, networks still generalize robustly; test accuracy remains high despite high clustering values. Flatness, however, is unaffected. In contrast, regularizing against flatness, i.e., encouraging sharp solutions, reliably delays generalization. These findings point to flatness as a more fundamental driver of generalization. To explain the frequent co-occurrence of NC and flatness, we provide a theoretical result: under classical NC assumptions, collapse implies relative flatness. This unifies previously observed correlations under a single geometric framework, but also shows that flatness can arise via other mechanisms. Neural collapse is thus a sufficient but non-exclusive path toward flatness. Finally, we revisit the theoretical framework of Petzka et al. (2021), which asserts that generalization requires not only flatness, but also representativeness, i.e., the alignment between learned features and the true data distribution. Our experiments confirm this dependency. Representativeness, when approximated, improves only as generalization sets in, highlighting that neither flatness nor collapse alone ensures generalization without meaningful feature alignment.
Together, these results place neural collapse in a new light. While it often accompanies generalization, it is not a prerequisite. Instead, it may serve as an inductive bias toward flatness under standard training dynamics. Relative flatness, by contrast, appears empirically necessary and theoretically justified, especially when labels are locally constant and features are representative—assumptions which appear to hold in our experimental settings.
@@
  • We prove that neural collapse implies relative flatness under classical assumptions;
  • We demonstrate that delayed generalization can be actively induced by manipulating flatness, even in real-world architectures.
@@ 2 Related Work
Despite substantial empirical success, the mechanisms driving generalization in deep neural networks remain poorly understood. Two pronounced geometric perspectives, neural collapse and loss surface flatness, have emerged as phenomena empirically correlated with generalization, but their interrelation and causal roles remain ambiguous. In this work, we leverage the properties of grokking to disentangle necessity from sufficiency and to assess which of these phenomena are fundamentally required for generalization.
@@
Neural collapse (NC) describes a geometric convergence during training a classifier where penultimate layer features from the same class collapse to a single mean, the means form a simplex equiangular tight frame, and classifier weights align accordingly (Papyan et al., 2020). This phenomenon has been widely observed at late training stages in overparameterized networks (Han et al., 2021; Graf et al., 2021; Rangamani and Banburski-Fahey, 2022; Wu and Papyan, 2024; Zhou et al., 2025) and has become a widely accepted indicator for generalization and robustness. Under idealistic assumption of an unconstrained features model, i.e., a model which optimizes features as free variables, NC has been shown to be the global optimal solution for various loss functions (Mixon et al., 2022; Zhou et al., 2022; Súkeník et al., 2023; Zhu et al., 2021). Similar results for compatibility of zero test error and collapse of variances (so only NC1 condition) were shown for mean-field networks (Wu and Mondelli, 2025). Practical applications have leveraged NC properties for downstream performance. For instance, Wu et al. (2025) use a feature separation loss based on NC to improve out-of-distribution detection. Similarly, Munn et al. (2024) show that lower geometric complexity in pre-trained representations fosters NC and leads to better few-shot generalization in transfer learning. Galanti et al. (2021) shows theoretically and empirically that NC is the reason for transfer learning to work better. These findings suggest that NC may be sufficient for generalization in certain regimes. However, the open question remains: is NC necessary outside of the unconstrained features models? For example, NC is a debated matter in the context of inbalanced learning, where it was observed that minority collapse can happen, i.e., the means of the minority classes merge thus preventing classification (Fang et al., 2021). At the same moment it is possible to induce NC by enforcing it through training, even in the unbalanced data setup (Yang et al., 2022). For deep linear networks it was also theoretically shown that global optimum exhibits NC, but even more: Under unbalanced data NC gets means proportional to the class size (Dang et al., 2023). Hui et al. (2022) argues for measuring NC on the test set for understanding generalization properties and demonstrates that test set NC is hurting generalization performance on the downstream tasks, i.e., representations with strong NC on the test set are bad for transfer learning. Also, large enough signal-to-noise ratio was observed to be an important criterion for NC to signify good generalization (Hong and Ling, 2024).
@@
The connection between loss surface flatness and generalization has long been suspected (Hochreiter and Schmidhuber, 1997), and more recently has been supported by empirical studies showing that flatness-based measures often outperform alternatives such as norm-, margin- or optimization-based metrics (Jiang et al., 2020; Keskar et al., 2017). However, classical flatness measures, typically based on the Hessian of the loss with respect to parameters, are known to be sensitive to reparameterizations that leave the function and generalization behavior unchanged (Dinh et al., 2017) and also very cumbersome to compute for large state-of-the-art models. To resolve this, Petzka et al. (2021) introduced a reparameterization-invariant measure of relative flatness, grounded in a theoretical framework that connects flatness of an individual layer to generalization under assumptions of representative data and locally constant labels. This formulation not only explains prior empirical correlations, but also provides theoretically grounded conditions under which flatness predicts generalization. Walter et al. (2024) further made this measure more practical by deriving closed-form computation formulas in standard classification settings. In our work, we empirically verify that relative flatness drops precisely when generalization emerges during grokking. Moreover, we demonstrate that minimizing relative flatness can actively induce generalization even in the absence of NC. In doing so, we confirm and extend the theoretical framework of Petzka et al. (2021), while disentangling flatness from NC.
@@
Grokking refers to the sudden emergence of generalization long after perfect training performance is achieved, as first described by Power et al. (2022). This phase transition, which is often induced by implicit or explicit regularization, offers a unique opportunity to isolate potential causes for the emergence of generalization. Recent work explains grokking through two-phase dynamics: an initial kernel regime focused on memorization, followed by feature learning that supports generalization (Mohamadi et al., 2023; Lyu et al., 2024; Kumar et al., 2024). Analytical studies reveal closed-form grokking solutions in modular arithmetic tasks (Gromov, 2024; Žunkovič and Ilievski, 2022), while mechanistic studies identify sparse and dense subnetworks competing during training (Nanda et al., 2023; Gouki et al., 2024). Most of these studies focus on shallow networks or toy tasks. However, recent work shows that grokking also arises in deeper architectures and real-world datasets (Murty et al., 2023; Humayun et al., 2024), suggesting greater relevance. We build on this foundation by using grokking as an experimental lens: By tracking the timing of flatness, NC, and generalization, we are able to distinguish which factors are causally connected and which are merely correlated.
@@ 3 Geometric Foundations of Generalization
This measure is invariant to neuron-wise rescaling and composition with orthogonal transformations, making it robust across different parameterizations. Under mild assumptions—namely, that the training data is representative and the labels are locally constant in feature space—relative flatness predicts generalization. Note that a small value of $\kappa^{\phi}_{\mathrm{Tr}}(\mathbf{w})$ is indicative of a flat solution and good generalization. In Section 5, we further show that NC implies low relative flatness, establishing a theoretical dependency between the two phenomena.
@@ 4 Neural Collapse and Flatness in Delayed Generalization
First, we study grokking on symbolic algorithmic tasks such as modular arithmetic (e.g., $x+y\mod p$), where networks are trained to predict the output of binary operations over abstract tokens. These tasks are small, fully observable, and require exact generalization beyond memorized samples, making them an ideal testbed. We follow the original setup of Power et al. (2022), training a 2 layer-transformer using the AdamW optimizer with a learning rate of $10^{-4}$ and weight decay of $1.0$. Each experiment runs for $10^{6}$ steps with a 50/50 train/validation split. All results are averaged over three seeds. Figure 1 shows the evolution of train and validation accuracy, as well as relative flatness (Def. 3.2), and neural collapse clustering (Def. 3.1) throughout training. We observe that both measures decrease sharply when the network start generalizing, however, NCC already decreases during the memorization phase. Relative flatness remains high during the memorization phase and decreases sharply near the point at which generalization begins. This temporal coincidence suggests that both NCC and relative flatness are correlated with generalization, but representations start exhibiting collapse behaviour before generalization is satisfying.
@@ 5 Neural Collapse is not Necessary
Neural collapse (NC) is often observed in models that generalize well, leading to the impression that it may play a functional role in generalization. However, correlation does not imply necessity. In this section, we demonstrate both empirically and theoretically that NC is not required for generalization. While collapse can accompany good generalization, it is neither sufficient nor necessary: it is one path that leads to flatness, and through flatness, to generalization.
@@
where $\mathcal{L}_{\text{CE}}$ is the standard cross-entropy loss and $NCC$ is the NC clustering measure (Def. 3.1). The NCC term penalizes tight clustering of penultimate-layer features by shrinking the distances between class means and increasing intra-class variance, thereby actively discouraging the geometric structure characteristic of NC. The training hyperparameter setups and additional experimental details are in Appendix B.
Figure 2 presents the results. As shown in panel (a), both training and validation accuracies remain unaffected by the regularizer. Panel (b) shows that relative flatness, our proxy for generalization, also remains stable. However, panel (c) reveals that NC is effectively suppressed: after a brief drop in the NCC measure, it steadily increases throughout training. This provides clear evidence that generalization can occur in the complete absence of NC, and that collapse is not a necessary condition for generalization. In Figure 3 we confirm that the NCC measure correctly captures cluster angles as well. Under neural collapse, angles are close to the optimal 10-simplex, yet when we regularize with the NCC measure, angles deviate strongly from this optimum.
Why, then, does neural collapse often accompany generalization? We argue that collapse is one geometric pathway to flat solutions, which are more directly linked to generalization. In fact, under mild assumptions, NC in the penultimate layer implies relative flatness [^1] Related works often rely on simplifying assumptions, e.g., one-layer behavior or layer peeling (Mixon et al., 2022; Fang et al., 2021), yet our analysis does not require such approximations.. This relationship is formalized below.
We assume that the classical neural collapse (Papyan et al., 2020) in the limit holds for a network function $f$. That is, the following four conditions apply to the neural network $f(x)=\text{softmax}(w\phi(x)+b)$ with $\phi(x)$ as the representation of the penultimate layer.
@@
    $\text{argmax}_{j}w_{j}^{\top},h+b_{j}=\text{argmin}_{j}\|h-\mu_{j}\|$.
@@
Let $f(x)=\text{softmax}(w\phi(x)+b)$ be a neural network with softmax output and trained with cross-entropy loss, where $w\in\mathbb{R}^{C\times d}$ denotes the final-layer weight matrix classifying into $k$ classes and $\phi(x)$ is the penultimate-layer representation. Assume that the classical neural collapse holds in the limit as above. Then relative flatness $\kappa_{\phi}(w)$ is bounded by:
@@
Taken together, our results demonstrate that, while NC can facilitate generalization by inducing flatness, it is not a prerequisite. Flatness remains the more fundamental quantity, and its relevance critically depends on the representativeness of the learned feature space. It remains to investigate whether flatness is a sufficient condition or whether it could be necessary.

### v2

@@ Abstract
Neural collapse, i.e., the emergence of highly symmetric, class-wise clustered representations, is frequently observed in deep networks and is often assumed to reflect or enable generalization. In parallel, flatness of the loss landscape has been theoretically and empirically linked to generalization. Yet, the causal role of either phenomenon remains unclear: Are they prerequisites for generalization, or merely by-products of training dynamics? We disentangle these questions using grokking, a training regime in which memorization precedes generalization, allowing us to temporally separate generalization from training dynamics and we find that while both neural collapse and relative flatness emerge near the onset of generalization, only flatness consistently predicts it. Models encouraged to collapse or prevented from collapsing generalize equally well, whereas models regularized away from flat solutions exhibit delayed generalization, resembling grokking, even in architectures and datasets where it does not typically occur. Furthermore, we show theoretically that neural collapse leads to relative flatness under classical assumptions, explaining their empirical co-occurrence. Our results support the view that relative flatness is a potentially necessary and more fundamental property for generalization, and demonstrate how grokking can serve as a powerful probe for isolating its geometric underpinnings. [^1] Code implementation: https://github.com/TrustworthyMachineLearning-Lab/grokking_flatness.
@@ 1 Introduction
In this paper, we challenge the common conjecture that NC is necessary for generalization (Mixon et al., 2022; Zhou et al., 2022; Súkeník et al., 2023; Zhu et al., 2021). Our key idea is to leverage grokking as a unique observational window: since generalization emerges only after prolonged memorization, it enables a clean separation between training dynamics and generalization. This allows us to ask: do NC and flatness merely correlate with generalization, or do they contribute to it?
We measure both phenomena in the penultimate layer. For flatness, we adopt the relative flatness metric proposed by Petzka et al. (2021), which considers the Hessian trace normalized by the weight norm, a quantity theoretically and empirically shown to align with generalization. To make this computable in large state-of-the-art neural networks, we employ the alternative closed-form upper bound introduced by (Walter et al., 2024), which is valid in the penultimate layer under cross-entropy loss. For NC, we measure the class-wise clustering of penultimate-layer representations through empirical variance and mean properties, tracking its emergence throughout training. In grokking experiments, we find that flatness emerges alongside generalization, never before. NC, in contrast, appears earlier during memorization, suggesting that it can lead to, but is not neccessary for generalization (cf. Fig. 1).
@@
To isolate their functional roles, we perturb training dynamics. When we suppress collapse-style clustering, networks still generalize robustly; test accuracy remains high despite high clustering values. Flatness, however, is unaffected. In contrast, regularizing against flatness, i.e., encouraging sharp solutions, reliably delays generalization and induces grokking-like behavior even in settings where it does not typically arise, including ResNets on CIFAR-10, ViTs on ImageNet and various pre-trained language models on SST-5. These findings point to flatness as a more fundamental driver of generalization. To explain the frequent co-occurrence of NC and flatness, we provide a theoretical result: under classical NC assumptions, collapse leads to relative flatness. This unifies previously observed correlations under a single geometric framework, but also shows that flatness can arise via other mechanisms. NC is thus a sufficient but non-exclusive path toward flatness. Finally, we revisit the theoretical framework of Petzka et al. (2021), which asserts that generalization requires not only flatness, but also representativeness, i.e., the alignment between learned features and the true data distribution. Our experiments confirm this dependency. Representativeness, when approximated, improves only as generalization sets in, highlighting that neither flatness nor collapse alone ensures generalization without meaningful feature alignment.
Together, these results place NC in a new light. While it often accompanies generalization, it is not a prerequisite. Instead, it may serve as an inductive bias toward flatness under standard training dynamics. Relative flatness, by contrast, tends to be empirically necessary and theoretically justified, especially when labels are locally constant and features are representative—assumptions which appear to hold in our experimental settings.
@@
  • we prove that neural collapse leads to relative flatness under classical assumptions;
  • we demonstrate that grokking-like behavior can be actively induced by manipulating flatness, even in real-world architectures.
@@ 2 Related Work
Despite substantial empirical success, the mechanisms driving generalization in deep neural networks remain poorly understood. Two pronounced geometric perspectives, neural collapse and loss surface flatness, have emerged as phenomena empirically correlated with generalization, but their interrelation and causal roles remain ambiguous. In this work, we leverage the properties of grokking to disentangle necessity from sufficiency and to assess which of these phenomena are required for generalization.
@@
Neural collapse (NC) describes a geometric convergence during training a classifier where penultimate layer features from the same class collapse to a single mean, the means form a simplex equiangular tight frame, and classifier weights align accordingly (Papyan et al., 2020). This phenomenon has been widely observed at late training stages in overparameterized networks (Han et al., 2021; Graf et al., 2021; Rangamani and Banburski-Fahey, 2022; Wu and Papyan, 2024; Zhou et al., 2025) and has become a widely accepted indicator for generalization and robustness. Under idealistic assumption of an unconstrained features model, i.e., a model which optimizes features as free variables, NC has been shown to be the global optimal solution for various loss functions (Mixon et al., 2022; Zhou et al., 2022; Súkeník et al., 2023; Zhu et al., 2021). Similar results for compatibility of zero test error and collapse of variances (so only NC1 condition) were shown for mean-field networks (Wu and Mondelli, 2025). Practical applications have leveraged NC properties for downstream performance. For instance, Wu et al. (2025) use a feature separation loss based on NC to improve out-of-distribution detection. Similarly, Munn et al. (2024) show that lower geometric complexity in pre-trained representations fosters NC and leads to better few-shot generalization in transfer learning. Galanti et al. (2021) shows theoretically and empirically that NC is the reason for transfer learning to work better. These findings suggest that NC may be sufficient for generalization in certain regimes. However, the open question remains: is NC necessary outside of the unconstrained features models? For example, NC is a debated matter in the context of inbalanced learning, where it was observed that minority collapse can happen, i.e., the means of the minority classes merge, thus preventing classification (Fang et al., 2021). At the same moment it is possible to induce NC by enforcing it through training, even in the unbalanced data setup (Yang et al., 2022). For deep linear networks it was also theoretically shown that the global optimum exhibits NC, but even more: Under unbalanced data, NC gets means proportional to the class size (Dang et al., 2023). Hui et al. (2022) argues for measuring NC on the test set for understanding generalization properties and demonstrates that test set NC is hurting generalization performance on related downstream tasks, i.e., representations with strong NC on the test set are bad for transfer learning. Moreover, Hong and Ling (2025) show that a sufficiently high signal-to-noise ratio is a necessary condition for NC to correlate with good generalization, explaining why strong NC can otherwise coincide with poor transfer performance.
@@
The connection between loss surface flatness and generalization has long been suspected (Hochreiter and Schmidhuber, 1997), and more recently has been supported by empirical studies showing that flatness-based measures often outperform alternatives such as norm-, margin- or optimization-based metrics (Jiang et al., 2020; Keskar et al., 2017). However, classical flatness measures, typically based on the Hessian of the loss with respect to parameters, are known to be sensitive to reparameterizations that leave the function and generalization behavior unchanged (Dinh et al., 2017), and computationally expensive for large state-of-the-art models. To resolve this, Petzka et al. (2021) introduced a reparameterization-invariant measure of relative flatness, grounded in a theoretical framework that connects flatness of an individual layer to generalization under assumptions of representative data and locally constant labels. This formulation not only explains prior empirical correlations, but also provides theoretically grounded conditions under which flatness predicts generalization. Walter et al. (2024) further made this measure more practical by deriving closed-form computation formulas in standard classification settings. In our work, we empirically verify that relative flatness drops precisely when generalization emerges during grokking. Moreover, we demonstrate that promoting relative flatness can actively induce generalization even in the absence of NC. In doing so, we confirm and extend the theoretical framework of Petzka et al. (2021), while disentangling flatness from NC.
@@
Grokking refers to the sudden emergence of generalization long after perfect training performance is achieved, as first described by Power et al. (2022). This phase transition, which is often induced by implicit or explicit regularization, offers a unique opportunity to isolate potential causes for the emergence of generalization. Recent work explains grokking through two-phase dynamics: an initial kernel regime focused on memorization, followed by feature learning that supports generalization (Mohamadi et al., 2023; Lyu et al., 2024; Kumar et al., 2024). Analytical studies reveal closed-form grokking solutions in modular arithmetic tasks (Gromov, 2023; Žunkovič and Ilievski, 2024), while mechanistic studies identify sparse and dense subnetworks competing during training (Nanda et al., 2023; Gouki et al., 2023). Most of these studies focus on shallow networks or toy tasks. However, recent work shows that grokking also arises in deeper architectures and real-world datasets (Murty et al., 2023; Humayun et al., 2024), suggesting greater relevance. We build on this foundation by using grokking as an experimental lens: By tracking the timing of flatness, NC, and generalization, we are able to distinguish which factors are causally connected and which are merely correlated.
@@ 3 Geometric Foundations of Generalization
This measure is invariant to neuron-wise rescaling and composition with orthogonal transformations, making it robust across different parameterizations. Under mild assumptions—namely, that the training data is representative and the labels are locally constant in feature space—relative flatness predicts generalization. Note that a small value of $\kappa^{\phi}_{\mathrm{Tr}}(\mathbf{w})$ is indicative of a flat solution and good generalization. In Section 5, we further show that NC leads to low relative flatness, establishing a theoretical dependency between the two phenomena.
@@ 4 Neural Collapse and Flatness in Delayed Generalization
First, we study grokking on symbolic algorithmic tasks such as modular arithmetic (e.g., $x+y\mod p$), where networks are trained to predict the output of binary operations over abstract tokens. These tasks are small, fully observable, and require exact generalization beyond memorized samples, making them an ideal testbed. We follow the original setup of Power et al. (2022), training a 2 layer-transformer using the AdamW optimizer with a learning rate of $10^{-4}$ and weight decay of $1.0$. Each experiment runs for $10^{6}$ steps with a 50/50 train/validation split. All results are averaged over three seeds. Figure 1 shows the evolution of train and validation accuracy, as well as relative flatness (Def. 3.2), and neural collapse clustering (Def. 3.1) throughout training. We observe that both measures decrease sharply when the network starts generalizing, however, NCC already decreases during the memorization phase. Relative flatness remains high during the memorization phase and decreases sharply near the point at which generalization begins. This temporal coincidence suggests that both NCC and relative flatness are correlated with generalization, but representations start exhibiting collapse behaviour before generalization has emerged.
@@ 5 Neural Collapse is not Necessary
Neural collapse (NC) is often observed in models that generalize well, leading to the impression that it may play a functional role in generalization. However, correlation does not imply necessity. In this section, we demonstrate both empirically and theoretically that NC is not required for generalization. While collapse can accompany good generalization, it is not necessary: It is one path that leads to flatness, and through flatness, to generalization.
@@
where $\mathcal{L}_{\text{CE}}$ is the standard cross-entropy loss and $NCC$ is the NC clustering measure (Def. 3.1). The NCC term penalizes tight clustering of penultimate-layer features by shrinking the distances between class means and increasing intra-class variance, thereby actively discouraging the geometric structure characteristic of NC. The training hyperparameter setups, additional experimental details are provided in Appendix B.
We show that generalization can occur without NC in Figure 2. As shown in panel (a), both training and validation accuracies remain unaffected by the regularizer. Panel (b) shows that relative flatness, our proxy for generalization, also remains stable. However, panel (c) reveals that NC is effectively suppressed: after a brief drop in the NCC measure, it steadily increases throughout training. This provides clear evidence that generalization can occur in the complete absence of NC, and that collapse is not a necessary condition for generalization. In Figure 3 we confirm that the NCC measure correctly captures cluster angles as well. Under neural collapse, angles are close to the optimal 10-simplex, yet when we regularize with the NCC measure, angles deviate strongly from this optimum.
Why, then, does neural collapse often accompany generalization? We argue that collapse is one geometric pathway to flat solutions, which are more directly linked to generalization. In fact, under mild assumptions, NC in the penultimate layer leads to relative flatness [^2] Related works often rely on simplifying assumptions, e.g., one-layer behavior or layer peeling (Mixon et al., 2022; Fang et al., 2021), yet our analysis does not require such approximations.. This relationship is formalized below.
We assume that classical neural collapse (Papyan et al., 2020) in the limit holds for a network function $f$. That is, the following four conditions apply to the neural network $f(x)=\text{softmax}(w\phi(x)+b)$ with $\phi(x)$ as the representation of the penultimate layer.
@@
    The classifier weights align with class means, i.e., $w_{j}=\lambda(\mu_{j}-\mu_{g})$ for all $j$ and some $\lambda>0$.
@@
    $\text{argmax}_{j}w_{j}^{\top}h+b_{j}=\text{argmin}_{j}\|h-\mu_{j}\|$.
Remark 5.2.
For the proof of Proposition 5.3, we additionally assume that $\|\mu_{c}\|\leq M$ for each $c$. This entails that the center of the ETF is not diverging. This mild technical condition does not alter the essence of the neural collapse assumptions.
@@
Let $f(x)=\text{softmax}(w\phi(x)+b)$ be a neural network with softmax output and trained with cross-entropy loss, where $w\in\mathbb{R}^{k\times d}$ denotes the final-layer weight matrix classifying into $k$ classes and $\phi(x)$ is the penultimate-layer representation. Assume that the classical neural collapse holds in the limit as above. Then relative flatness $\kappa_{\phi}(w)$ is bounded by:
@@
Taken together, our results demonstrate that, while NC can facilitate generalization by inducing flatness, it is not a prerequisite. Flatness remains the more fundamental quantity, and its relevance critically depends on the representativeness of the learned feature space. It remains to investigate whether flatness is a sufficient condition and whether it could be necessary.